find linearly independent function that are annihilated by the given differential operator D⁵,(D²+4D)
1
Expert's answer
2020-12-02T10:40:59-0500
"Solution"

The linearly differential operator "D^5(D^2+4D)" which is "D^6(D+4)" in the form "D^{n+1}(D +\alpha )". Here "n=5" and hence it can annihilate the independent functions that are in the form "x^{n \le5}" and "e^{-\alpha x}", then we can have the functions being annihilated by the given differential operator as "x^0, x^1, x^2, x^3, x^4,x^5" and "e^{-4x}".

"\therefore 1, x, x^2, x^3, x^4,x^5\ and\ e^{-4x}"


Where "n=0,1,2,3,4,5\ and\ \alpha=4"